# Identifying Best Candidates for Busbar Splitting

### v1

@@ 4.3 Validation Results
The rightmost column of Table 4 show the cost decrease for the identified busbars. As for the 118-bus case, selecting busbars with high $\phi$ values and larger # of connected branches/elements leads to cost decreases for several busbars, namely 693, 470, 649, 171 and 23. At the same time, the 7 busbars selected using the # of branches/elements and $\phi$ values above the mean-$\phi$ lead to cost decreases, too.
To validate whether we identified the “best” busbars for splitting, we validate our results by splitting every busbar in the system and checking the cost reduction. Fig. 4 shows the results obtained by splitting one busbar at a time for all busbars in the 793-bus system. We observe that the proposed metrics identify 3 out of the 4 busbars leading to the highest reductions in generation costs, as well as several other buses among the ones with high cost reductions. Busbar 444, which leads to the third-best cost decrease if split, has only two connected branches, both utilized above 80% of their rated capacity (one at 100%). As with most busbars in the bottom part of Table 3, not reconnecting a branch to the split busbar is leading to 444’s reduction in costs, suggesting that line switching rather than BuS is what achieves the cost reduction.
Note that for this test case, splitting one busbar per time took 68102 s (19.14 h), 67639 s for the LPAC-BuS, and 463.3 s for the AC-FC, while the total computational time for the selected buses is considerably lower, with 3523.3 s (0.98 h).
Figure 4: Cost decrease of the AC-OPF feasibility check for the optimized topology of the selected busbars compared to the results obtained by splitting one busbar at a time for case 793-bus. Most of the busbars selected with the identified metrics do lead to a reduction in generation costs. The busbars with the highest cost decrease tend to have a sum of the absolute differences in LMPs over the branches connected to it ($\phi$) higher than the mean $\phi$ value.
4.3.2 3374-bus test case
We use similar metrics as the previous case to identify relevant candidates for BuS. On the top part of Table 5, the first 13 busbars are selected by considering the 20 busbars with the highest $\phi$-values, and including the only ones with the highest $\phi$ and # of connected elements/branches. Note that busbars 679, 611 and 665 are selected even if they have a low number of branches, as their $\phi$-values are high. In particular, busbars 679 and 611 have $\xi=1$ and only two branches, as busbar 444 in the previous case, possibly leading to potential for line switching. The bottom part of Table 5 shows 7 busbars selected according to their # of connected elements/branches and their above-average $\phi$ value. The mean $\phi$ value in this case is 483.55 [\$/MWh].
Table 5: Identification of 20 candidates for busbar splitting in the 3374-bus test case with the proposed metrics, and related cost decrease.
@@
Table 5 and Fig. 5 show how 8 of the selected busbar leads to reductions in generation costs compared to the AC-OPF. Also in this case, part of the selected busbars, i.e. 498, 665, 439, and 670, figure among the ones leading to the best cost decrease. Different from case 793-bus, the busbars in the bottom part of Table 5 do not lead to any cost decrease even though they have many connected lines and elements. This fact can be related to the lower cost reductions compared to the previous cases, showing that large cost reductions are not achievable in this specific test case, and it is therefore more complex to identify relevant busbars for BuS. For example, busbar 10001 leads to the biggest cost reduction even though it has 4 non-congested branches connected to it and a low $\phi$, making it difficult to be identified as a good candidate from the OPF results. Nevertheless, Fig. 5 still confirms the validity of the proposed metrics in identifying relevant busbars for BuS.
Figure 5: Cost decrease of the AC-OPF feasibility check for the optimized topology of the selected busbars compared to the results obtained by splitting one busbar at a time for case 3374-bus. The proposed metrics identify several busbars leading to the highest reductions in generation costs.
Finally, optimizing the selected buses took 13582.7 s (3.77 h), or an average of 11.32 min per bus. Therefore, the proposed metrics select relevant busbars in a substantially lower time than if we had to split one busbar at a time, which took 374.98 h (15.62 days).
5 Conclusion and future work
@@
Specifically, the results show that our newly proposed $\phi$-metric, defined as the sum of the absolute differences in LMPs over the branches connected to a busbar ($\phi$), in combination with the # of elements connected to the busbar can help identify the busbars which achieve the highest cost reductions. Further, the presence of binding voltage angle and magnitude limits ($\zeta$-metric) indicate that busbar splitting will likely not reduce cost. Some of the buses that were not identified by our metrics achieved cost reductions through transmission line switching rather than busbar splitting.
Given the flexibility of the proposed grid topology optimization model in including more than one busbar, future work will include testing combinations of busbars to achieve further reductions in generation costs. Moreoverx, the proposed metrics can be used in grid planning side investigate the possibility of retrofitting existing substations to effective topological actions when congestion takes place in the grid or new substations or RES capacity are to be installed in the grid.

### v2

@@ IV-B Validation phase results
The rightmost column of Table V shows the cost decrease for the identified busbars. As for the 118-bus case, selecting busbars with high $\phi$ values and larger # of connected branches/elements leads to cost decreases for several busbars, namely 693, 470, 649, 171 and 23. At the same time, the 7 busbars selected using the # of branches/elements and $\phi$ values above the mean-$\phi$ lead to cost reductions as well.
To validate whether we identified the “best” busbars for splitting, we validate our results by splitting every busbarin the system and checking the cost reduction. Fig. 4 shows the results obtained by splitting one busbar at a time for all busbars in the 793-bus system. We observe that the proposed metrics identify 3 of the 4 busbars that lead to the largest reductions in generation costs, as well as several other buses with high cost reductions. We note that busbar 444, which yields the third-best cost reduction if split, has only two connected branches, both of which are utilized above 80% of their rated capacity (one at 100%). As with most busbars in the bottom part of the previous Table IV, not reconnecting a branch to the split busbar is leading to 444’s reduction in costs, suggesting that line switching rather than BuS is what achieves the cost reduction.
For this test case, splitting one busbar per time took 68102 s (19.14 h), 67639 s for the LPAC-BuS and 463.3 s for the AC-FC, while the total computational time to optimize only the the selected buses is considerably lower, with 3523.3 s (0.98 h).
Figure 4: Cost decrease of the AC-OPF feasibility check for the optimized topology of the selected busbars compared to the results obtained by splitting one busbar at a time for the 793-bus test case. Most of the busbars selected with the identified metrics do lead to a reduction in generation costs. The busbars with the highest cost decreases tend to have a sum of the absolute differences in LMPs across the branches connected to them ($\phi$) higher than the mean $\phi$ value.
IV-B2 3374-bus test case
Finally, we use the same metrics as in the previous case to identify relevant candidates for BuS for the 3374-bus test case, too. On the top part of Table VI, the first 13 busbars are selected by considering the 20 busbars with the highest $\phi$-values, and including the only ones with the highest $\phi$ and # of connected elements/branches. Note that busbars 679, 611 and 665 are selected even if they have a low number of branches, as their $\phi$-values are high. In particular, busbars 679 and 611 have $\xi=1$ and only two branches, as busbar 444 in the previous case, possibly leading to potential for line switching. The bottom part of Table VI shows 7 busbars selected based on their # of connected elements/branches and their above-average $\phi$ values. The mean $\phi$ value in this case is 483.55 [\$/MWh].
TABLE VI: Identification of 20 candidates for busbar splitting in the 3374-bus test case with the proposed metrics, and related cost decrease.
@@
Table VI and Fig. 5 show that 8 of the selected busbar leads result in reductions in generation costs compared to the AC-OPF. Also in this case, part of the selected busbars, i.e., 498, 665, 439, and 670, are among those that lead to the largest cost decrease. Unlike case 793-bus, the busbars in the bottom part of Table VI do not lead to any cost reduction, even though they have several lines and elements connected to them. This result can be attributed to the lower cost reductions relative to previous test cases, indicating that large cost reductions are not achievable in this specific test case, and that identifying relevant busbars for BuS is therefore more challenging. For example, busbar 10001 yields the largest cost reduction, even though it has 4 non-congested branches connected to it and a low $\phi$, making it difficult to identify as a good candidate from the OPF results. In addition, out of the busbars with high $\phi$-values and low number of branches (679, 611 and 665), only 665 leads to a cost decrease if split. We leave further refinements of the metrics to efficiently select busbars with similar characteristics for future work. Nevertheless, Fig. 5 still confirms the validity of the proposed metrics in identifying relevant busbars for BuS, with several selected busbars being among the ones leading to the highest cost reductions compared to the AC-OPF simulation.
Figure 5: Cost decrease of the AC-OPF feasibility check for the optimized topology of the selected busbars compared to the results obtained by splitting one busbar at a time for the 3374-bus test case. The proposed metrics identify several busbars leading to the highest reductions in generation costs.
Finally, optimizing the selected buses took 13582.7 s (3.77 h), or an average of 11.32 min per bus. Therefore, the proposed metrics select relevant busbars substantially faster than splitting one busbar at a time, which took 374.98 h (15.62 days).
V Conclusion and future work
@@
Specifically, the results show that our newly proposed $\phi$-metric, defined as the sum of the absolute differences in LMPs over the branches connected to a busbar ($\phi$), in combination with the # of elements connected to the busbar, can help identify the busbars that achieve the highest cost reductions. Further, the presence of binding voltage angle and magnitude limits ($\zeta$-metric) indicates that busbar splitting will likely not reduce cost. Some of the buses that were not identified by our metrics achieved cost reductions through transmission line switching rather than busbar splitting.
Given the flexibility of the proposed grid topology optimization model in including more than one busbar, future work will include testing combinations of busbars to achieve further reductions in generation costs. Moreover, the proposed metrics can be used in grid planning side to investigate the possibility of retrofitting existing substations to effective topological actions when congestion takes place in the grid, or new substations or RES capacity are installed in the grid.
Acknowledgement
This work is supported by the Belgian Energy Transition Fund, FOD Economie, project DIRECTIONS, and a Research Foundation – Flanders (FWO) travel grant for Giacomo Bastianel’s research visit to UW-Madison.